OpenStudy (anonymous):

Calculate the solubility of AgCl in a solution containing 0.1M MgCl2 and 0.2M AgNO3 I dont understand why here the reaction moves to reactant When doing the ICE table, the "C" is -x for Ag+ and Cl-

OpenStudy (anonymous):

AgCl-> Ag+ + Cl- Now There is already Ag+ and Cl- in the medium as MgCl2 and AgNo3 are strong electrolytes. Suppose new solubility of AgCl is s AgCl-> Ag+ + Cl- s+0.2 s+0.2 Ag+ is now s+0.2(0.2 comes from AgNo3) now solve it as (s+0.2)^2=Ksp

OpenStudy (anonymous):

Reaction moves to rectants side according to Le Chateliers principle.

OpenStudy (anonymous):

so do you determine if it moves to reactant side or product side by the Ksp value? I've learnt that Ksp < 0 : moves to reactant Ksp > 0 : moves to product

OpenStudy (anonymous):

No You are talking about ionic product Ksp is a constant irrespective of the conditions,Anyway if you solve s^2 + Ksp and then find solubility in pure water,you'll find s will be more for pure water than here in this solution. Ksp never changes.

OpenStudy (anonymous):

Ksp is always a positive constant.

OpenStudy (anonymous):

but this question given that Ksp AgCl = 1.8 x 10^-10 so it is not related on the reaction shifting reactant or to products?

OpenStudy (anonymous):

Ksp has nothing to do with it ,it is the solubility s that you should find. Let us take AgCl pure water AgCl-> Ag+ + Cl- s s s^2=1.8*10^-10= approx 1.4*10^-5 now lets take AgCl in the given solution. AgCl-> Ag+ + Cl- s+0.2 s+0.2 (s+0.2)^2=1.8*10^-10. Now if you calculate s you get a negative quantity, That basically means the solution is beyond saturated and AgCl starts to precipiate, If AgCl starts to precipitate then reaction is obviously moving to reactants side. I think what you are trying to ask is about ionic product....

OpenStudy (anonymous):

Ionic product= (Ag+)* (Cl-) In any sparingly soluble salt if ionic product is greater than Ksp it means that reaction will proceed in backward direction with AgCl precipiating until the concentration of Ag+ and Cl- has been reduced till ionic product is equal to Ksp. If ionic priduct is less than Ksp then the reaction will proceed forward with more AgCl disocciating to give Ag+ and Cl- Till again the ionic product becomes equal to Ksp.

OpenStudy (anonymous):

i didnt learn about the pure water stuff yet :( and yea, i think i'm talking about the ionic product too my teacher didnt talk much about it we need to find Q and compare it with Ksp if Q is greater than K, then precipitate forms

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

but then since i didnt learn about the pure water stuff, so i cant use ur way to determine if ppt forms or not if i do 1.8 x 10^-10 = (0.2 -x)^2 x = 0.1999 and 1.8x 10 ^-10 = (0.2+x)^2 x= - 0.1999 is there another way to detminer if ppt forms other then compare it to water?

OpenStudy (anonymous):

and thanks for ur help!

OpenStudy (anonymous):

Well the ionic product method is a better method just find the concentrations of the products and multiply them and compare it to Ksp, Finding s is an uneccessary step, I just told you that since you wanted solubility.

OpenStudy (anonymous):

ohhhh i get it now! so 0.2 x 0.2 = 0.04 and is larger than Ksp, that means precipitate forms!! oh thanks thanks thanks!

OpenStudy (preetha):

Great work. Thanks Shanvee for your explanation and thanks LovekBlue for seeking to understand!